OpenStudy (anonymous):

The diagonals of a square each measure 7 feet. Find the are of the square

OpenStudy (anonymous):

@Loujoelou

OpenStudy (ajprincess):

 Use pythagoras theorem \(a^2+a^2=7^2\) \(2a^2=49\) Area=\(a^2\) find \(a^2\) from \(2a^2=49\)